OpenStudy (anonymous):

Use laplace transforms to solve y''(t)+4y(t)=theta(t) with y(0)=1 y'(0)=0 where theta(t)= 4t, when t is in [0,1] and theta(t)=4 when t>1

OpenStudy (sidsiddhartha):

\[L[y'(t)=sY(s)-Y(0)\\L[y''(t)]=s^2Y(s)-sY(s)-Y'(0)\]

ganeshie8 (ganeshie8):

start by taking laplace transform through out and solve Y

OpenStudy (sidsiddhartha):

yes the above property shows u the laplace of y'(t) andy''(t)

OpenStudy (sidsiddhartha):

now just put them in the equation

OpenStudy (sidsiddhartha):

\[y''(t)+4y(t)=4t\\s^2Y(s)-sY(s)-Y'(0)+4Y(s)=\frac{ 4 }{ s^2 }\] now put the initial conditions

OpenStudy (anonymous):

Yeah I got that far but I got stuck with the theta. Does this mean I need to solve the same equation two different ways, for 4t and 4?

OpenStudy (anonymous):

to solve will I need to take the inverse once I find the laplace transform?

ganeshie8 (ganeshie8):

right hand side is an unknown function right ?

OpenStudy (anonymous):

right hand side is a function theta that can be either 4t or 4, depending on what t is

ganeshie8 (ganeshie8):

okay i see

OpenStudy (anonymous):

that's the confusing part haha

ganeshie8 (ganeshie8):

\[\large \begin{align} y''(t)+4y(t)&=\theta (t)\\~\\ s^2Y(s)-sY(s)-Y'(0)+4Y(s)&=\Theta(s)\\~\\ Y(s) &= \dfrac{1}{s^2-s+4}\Theta(s)\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

next you can try convolution

OpenStudy (anonymous):

woah....how did you get that above...? i thought you would split it into two equations and take separate laplace transforms

ganeshie8 (ganeshie8):

you will need to work two separate eqns for sure, above setup just postpones the splitting part till evaluating the integral :)

OpenStudy (anonymous):

ooooooooooooh gotcha!

ganeshie8 (ganeshie8):

\[\large \mathcal{L^{-1}}\left(\dfrac{1}{s^2-s+4} \right) = \frac{2}{\sqrt{15}}e^{t/2}\sin(t\sqrt{15}/2)\]

OpenStudy (anonymous):

so there is only one solution....not two separate solutions? I'm sorry that still just throws me for a loop lol

ganeshie8 (ganeshie8):

\[\large \begin{align} &y''(t)+4y(t)=\theta (t)\\~\\ &s^2Y(s)-sY(s)-Y'(0)+4Y(s)=\Theta(s)\\~\\ &Y(s) = \dfrac{1}{s^2-s+4}\Theta(s)\\~\\ &y(t)=\frac{2}{\sqrt{15}}\int\limits_0^t e^{u/2}\sin(u\sqrt{15}/2)~\theta(t-u)~\mathrm{d}u \end{align}\]

ganeshie8 (ganeshie8):

plugin the value of theta function and evaluate two times, somehow i feel not using convlution gives a smooth solution now.. :O

OpenStudy (anonymous):

oh Lord lol. So use convolution or no?

ganeshie8 (ganeshie8):

upto you

OpenStudy (anonymous):

ok lol. Thanks for your help!

ganeshie8 (ganeshie8):

can you work above messy integral ?

OpenStudy (anonymous):

ummmm.....honestly not really lol. But I have alot of free time tomorrow so I was gonna muck through it till I figured it out

ganeshie8 (ganeshie8):

sounds good, let me give it to wolfram quick and see what it spits out

OpenStudy (anonymous):

okie doke

OpenStudy (anonymous):

ok I'll try lol. If i run into problems tomorrow I will let you know!